# Constructing a 3D Town from a Single Image

### v1

@@ 4.1 Experimental Setup
To the best of our knowledge, there is no established benchmark for 3D outdoor scene mesh generation from single images. Therefore, we construct a custom test set by prompting GPT-4o [17] to generate 100 diverse top-down scene images in a variety of styles, such as “snow village,” “desert town,” and more. The generation details can be found in Appendix C.
Metrics  In the absence of ground-truth meshes, we measure performance through pairwise comparisons between models. For every reference image, two generated scenes are evaluated on three criteria: geometry quality, layout coherence, and texture coherence. Geometry quality assesses which mesh contains more detailed and fine-grained structures closer to the image evidence. Layout coherence evaluates whether the generated mesh has the same layout as the reference image. Texture coherence measures how well the generated mesh texture aligns with the corresponding parts of the reference image. Each pair is scored by two independent Amazon Mechanical Turk annotators, yielding a human-preference win rate. We additionally obtain a GPT-4o-based weighted win rate: GPT-4o is shown the same pair, returns token probabilities for alternatives “A” and “B,” and we treat $P(\text{A})$ as a soft vote when our method is option A (and analogously for B). Weighted win rate is computed as $P(\text{win})$ if our model wins, and $1-P(\text{lose})$ otherwise, then averaged across all pairs. All scenes are rendered to RGB images with identical Blender settings; further GPT prompt details appear in Appendix C.
Baselines  We compare 3DTown with three state-of-the-art image-to-3D generation methods: (1) Trellis [59] is a structured latent model that uses Rectified Flow Transformers to generate 3D assets from text or image prompts, supporting outputs such as radiance fields, 3D Gaussians, and meshes. (2) Hunyuan3D-2 [52] adopts a two-stage pipeline that separately synthesizes geometry and texture using a flow-based diffusion transformer, enabling high-resolution asset generation. (3) TripoSG [24] is a high-fidelity 3D shape synthesis model that employs large-scale rectified flow transformers and hybrid geometric supervision to generate detailed 3D meshes from single images, demonstrating strong generalization across diverse styles. They use MV-Adapter [16] for texture generation. All models are evaluated in a zero-shot setting using official pretrained checkpoints. To ensure fair comparison, background removal is disabled and the full input image is encoded for all methods.
@@ 4.2 Main Results
Table 1: Quantitative comparisons of 3DTown and baselines. We report human preference win rates and GPT-4o-based weighted win rates (%) across geometry, layout, and texture quality. 3DTown consistently outperforms all baselines by large margins in both evaluations.
@@
Trellis [59] 31.50% 30.00% 33.00% 17.58% 19.04% 14.75%
3DTown (Ours) 68.50% 70.00% 67.00% 82.42% 80.96% 85.25%
Hunyuan3D-2 [52] 33.50% 33.50% 31.50% 11.66% 12.13%   7.67%
3DTown (Ours) 66.50% 66.50% 88.34% 88.34% 87.87% 92.33%
TripoSG [24] 22.50% 23.00% 26.00% 24.60% 22.34% 22.79%
3DTown (Ours) 77.50% 77.00% 74.00% 75.40% 77.66% 77.21%
Figure 3: Qualitative comparisons between 3DTown and baselines. Given a single top-down image (left column), we compare 3D scene outputs generated by 3DTown, Trellis [59], Hunyuan3D-2 [52], and TripoSG [24]. 3DTown consistently produces globally coherent scenes with fine-grained geometry, accurate object layouts, and realistic textures across a variety of styles and environments. In contrast, Trellis often produces oversimplified geometry, Hunyuan3D-2 suffers from structural inconsistencies and domain mismatch, and TripoSG exhibits repetition artifacts and layout drift.
Table 1 and Figure 3 present the main quantitative and qualitative comparisons between 3DTown and baseline methods. The results clearly demonstrate that 3DTown consistently outperforms Trellis [59], Hunyuan3D-2 [52], and TripoSG [24] across geometry, layout, and texture quality, as evaluated by both human annotators and GPT-4o. These improvements can be attributed to our region-based design and spatially guided generation strategy, which together promote better alignment between image features and 3D content, while preserving scene-wide consistency.
@@
3DTown ’s region-wise decomposition enables each latent block to align with a localized image crop, narrowing the domain gap between object-centric training and scene-level inference. This design results in significant improvements in texture fidelity, with GPT-4o assigning 3DTown a texture win rate of 92.3% versus 7.7% for Hunyuan3D-2. Furthermore, the upscale of resolution allows for more fine-grained structure representation, which reflects on the improvements in geometry quality, where human preferences increase by 37 percentage points over Trellis (68.5% vs. 31.5%) and 55 points over TripoSG (77.5% vs. 22.5%).
Meanwhile, the use of spatial priors and masked 3D inpainting stabilizes region layout and inter-region transitions, and enforces global layout constraints, significantly reducing spatial artifacts and disjoint structures. This is reflected in layout coherence scores: 3DTown achieves 70.0% in human preference compared to 30.0% for Trellis, and reaches 87.9% in GPT-4o evaluation against only 12.1% for Hunyuan3D-2.
@@
Qualitatively, 3DTown produces scene assets with clear structure, consistent layout, and realistic surface details that closely match the reference top-down image. In contrast, Trellis often generates overly centralized, low-resolution structures and lacks peripheral detail. Hunyuan3D-2 exhibits notable issues with layout distortion and geometry hallucinations despite acceptable textures in isolated parts. TripoSG maintains some compositional structure but frequently introduces repeated objects and ignores the layout evidence within the reference image. 3DTown ’s region-wise generation and spatial inpainting pipeline helps it avoid these artifacts while maintaining both global coherence and local fidelity.
These findings confirm that spatial decomposition and prior-guided inpainting are effective principles for lifting single-view image inputs into coherent, high-quality 3D scenes. Additional qualitative comparisons are available in Figure 5 in Appendix A.
@@ 4.3 Ablation Study
Table 2: Ablation Study Results. Win rates for geometry, layout, and texture show that removing region-based generation or landmark conditioning degrades performance, highlighting the importance of both components in 3DTown.
@@
3DTown (w/o regions) 20.00% 17.00% 13.50% 6.11% 8.48% 6.08%
3DTown 80.00% 83.00% 86.50% 92.89% 91.52% 92.92%
3DTown (w/o landmarks) 41.50% 46.00% 42.00% 36.90% 44.21% 38.26%
3DTown 59.50% 54.00% 58.00% 64.10% 56.79% 61.74%
Figure 4: Qualitative Ablation Results. Left: Reference image. Middle: 3DTown without landmark conditioning. Right: 3DTown without region-based generation. Landmark conditioning ensures consistency for foreground objects, especially for objects across regions, while region-based generation preserves overall detail and coherence.
We conduct ablation studies to evaluate the contributions of key components in 3DTown: the region-based generation strategy and the use of pre-generated landmarks. Both ablation studies still apply the spatial-aware 3D completion. Quantitative results are shown in Table 2, and qualitative comparisons are illustrated in Figure 4.
@@
Here, we disable the landmark-aware initialization and rely only on monocular depth to construct the spatial prior. This results in a notable decrease in geometry and layout quality, particularly in regions containing large foreground structures like gates or towers. Because region-based generation processes the scene in spatial chunks, landmarks act as anchors to maintain object continuity across patch boundaries. Without them, the model is more likely to hallucinate disjoint or mismatched content between regions. Figure 4 (middle) shows a representative failure: the landmark structure in the right-side building is incorrectly completed and misaligned with its neighboring parts.
These findings demonstrate that both region-wise decomposition and landmark conditioning are essential for producing coherent, high-fidelity 3D scenes. Regions allow pretrained object generators to operate within their effective receptive fields, while landmarks preserve global anchors and prevent semantic drift during inpainting.
@@ 5 Conclusion
To address the challenge of generating high-quality, coherent 3D scenes from a single image, we proposed 3DTown, a training-free framework that synthesizes 3D scenes by decomposing them into overlapping regions and guiding generation with spatial priors. Central to our approach is the use of spatial-aware 3D completion, which enables localized object fidelity while maintaining global scene coherence through masked rectified flow. Empirical results show that 3DTown outperforms existing methods across geometry, texture, and layout quality. These findings highlight the potential of modular, spatially grounded generation for scalable 3D scene synthesis from minimal input.
@@ Appendix A More Results
Figure 5 presents additional qualitative results comparing 3DTown with Trellis [59], Hunyuan3D-2 [52], and TripoSG [24] across a diverse set of visual scenes. These examples further demonstrate the robustness and generality of our approach across different architectural styles, spatial layouts, and artistic domains.
3DTown consistently produces scene assets that are geometrically detailed, visually coherent, and well-aligned with the input reference images. In contrast, baseline models frequently suffer from artifacts such as repeated structures, layout collapse, or low-resolution textures. These comparisons further highlight the benefits of our region-based generation and spatial-aware inpainting pipeline in generating diverse 3D scenes from a single image without training.
@@ Appendix C Experiment Settings
To evaluate models’ performance under stylistically diverse conditions, we curated a human-verified synthesized image test set with 100 top-down views. Given an example image, we ask ChatGPT-o3 [36] to generate image prompts for top-down scene views with this requirments: 1280 × 720 resolution, quasi-orthographic three-quarter (“isometric”) camera, one hero landmark at the image centre, 10–20 surrounding buildings, daylight illumination, and “no far-away object”. Then, we used GPT-4o [17] to generate scene images according to the corresponding prompts.
@@
For GPT-based automatic evaluation, we ask the same questions as the human evaluation and prompt the model to directly return the answer with top-5 token log probability. Then, we extract the token probability of ‘A’ and ‘B‘ (0 if not included) as the answer weights.
@@ Appendix D Limitation
Although 3DTown delivers strong scene-level results, several challenges remain. The pretrained 3D generator we adopt is trained on single-object imagery; even after region decomposition, the underlying distribution mismatch can lead to patch-level hallucinations—for example, duplicated façades or unrealistic roof shapes. Future work could mitigate this via scene-level fine-tuning or domain adaptation.

### v2

@@ 4.1 Experimental Setup
To the best of our knowledge, there is no established benchmark for 3D outdoor scene mesh generation from single images. Therefore, we construct a custom test set by prompting GPT-4o (Hurst et al., 2024) to generate 100 diverse top-down scene images in a variety of styles, such as “snow village,” “desert town,” and more. The generation details can be found in Appendix C.
Metrics Without ground-truth meshes, we evaluate pairwise model comparisons per reference image on three criteria, geometry quality, layout coherence, and texture coherence, assessing detail fidelity, spatial arrangement, and texture–image alignment, respectively. Each pair is annotated by two AMT workers to yield a human-preference win rate. We also report a GPT-4o weighted win rate: given the same pair, GPT-4o outputs token probabilities for “A/B”; we use the probability of the chosen option as a soft vote and average over pairs. All scenes are rendered to RGB images with identical Blender settings; further GPT prompt details appear in Appendix C. Additionally, we report rendered-view image metrics between each method’s render and its reference image: CLIP similarity (openai/clip-vit-base-patch32, higher is better) and FID (Inception-V3, lower is better), averaged over the test set; these complement the human/GPT criteria by measuring image-level fidelity.
Baselines  We compare SceneFuse-3D with four image-to-3D generation methods: Trellis (Xiang et al., 2024), Hunyuan3D-2 (Team, 2025), TripoSG (Li et al., 2025), and LGM (Tang et al., 2024b). While Trellis, Hunyuan3D-2 and TripoSG represent the state-of-the-art end-to-end 3D transformer models, LGM represents the multi-view generation-based 3D generator. All models are evaluated in a zero-shot setting using official pretrained checkpoints. To ensure fair comparison, background removal is disabled, and the full input image is encoded for all methods. We also experimented with the progressive novel-view method WonderWorld (Yu et al., 2025). As it does not yield a consistent editable mesh, it is not included in our mesh-based quantitative tables; we provide qualitative comparisons and discussion in Appendix A.2.
@@ 4.2 Main Results
Table 1: Quantitative comparisons of SceneFuse-3D and baselines. We report human preference win rates and GPT-4o-based weighted win rates (%) across geometry, layout, and texture quality. SceneFuse-3D consistently outperforms all baselines by large margins in both evaluations.
@@
Trellis (Xiang et al., 2024) 31.50% 30.00% 33.00% 17.58% 19.04% 14.75%
SceneFuse-3D (Ours) 68.50% 70.00% 67.00% 82.42% 80.96% 85.25%
Hunyuan3D-2 (Team, 2025) 33.50% 33.50% 31.50% 11.66% 12.13%   7.67%
SceneFuse-3D (Ours) 66.50% 66.50% 88.34% 88.34% 87.87% 92.33%
TripoSG (Li et al., 2025) 22.50% 23.00% 26.00% 24.60% 22.34% 22.79%
SceneFuse-3D (Ours) 77.50% 77.00% 74.00% 75.40% 77.66% 77.21%
LGM (Tang et al., 2024b) 0.00% 0.00% 0.00% 0.60% 0.00% 0.00%
SceneFuse-3D (Ours) 100.00% 100.00% 100.00% 100.00% 100.00% 100.00%
Figure 3: Qualitative comparisons between SceneFuse-3D and baselines. Given a single top-down image (left column), we compare 3D scene outputs generated by SceneFuse-3D, Trellis (Xiang et al., 2024), Hunyuan3D-2 (Team, 2025), TripoSG (Li et al., 2025), and LGM (Tang et al., 2024b). SceneFuse-3D consistently produces globally coherent scenes with fine-grained geometry, accurate object layouts, and realistic textures across a variety of styles and environments. In contrast, Trellis often produces oversimplified geometry; Hunyuan3D-2 suffers from structural inconsistencies and domain mismatch; TripoSG exhibits repetition artifacts and layout drift; LGM cannot generate consistent multi-view scene images for 3D construction.
Table 12 and Figure 3 present the main quantitative and qualitative comparisons between SceneFuse-3D and baseline methods. The results clearly demonstrate that SceneFuse-3D consistently outperforms Trellis (Xiang et al., 2024), Hunyuan3D-2 (Team, 2025), TripoSG (Li et al., 2025), and LGM (Tang et al., 2024b) across geometry, layout, and texture quality, as evaluated by both human annotators and GPT-4o. These improvements can be attributed to our region-based design and spatially guided generation strategy, which together promote better alignment between image features and 3D content, while preserving scene-wide consistency.
@@
SceneFuse-3D ’s region-wise decomposition aligns each latent block with a localized image crop, reducing the domain gap between object-centric training and scene-level inference. This design boosts texture fidelity, with GPT-4o assigning a 92.3% win rate versus only 7.7% for Hunyuan3D-2. The upscaled resolution further enhances structural detail, reflected in geometry improvements of +37 points over Trellis (68.5% vs. 31.5%) and +55 points over TripoSG (77.5% vs. 22.5%). Spatial priors and masked 3D inpainting stabilize layouts and smooth inter-region transitions, yielding higher layout coherence (70.0% vs. 30.0% for Trellis in human study; 87.9% vs. 12.1% for Hunyuan3D-2 in GPT-4o evaluation). LGM, leveraging multi-view image generation, fails to provide consistent geometry in this setting and collapses to oversimplified outputs.
Rendered-view metrics in Table 2 support these findings. SceneFuse-3D achieves the highest CLIP similarity (0.8030), indicating stronger semantic and visual alignment with the input than all baselines. It also records the best FID (258.74), showing the closest distributional match in image structure, while Trellis follows at 288.23, and the others exceed 300. Together, these results confirm that SceneFuse-3D not only excels in human and GPT-4o preference studies but also in reference-based image metrics, underscoring its advantages in both semantic fidelity and distributional realism.
@@
Qualitatively, SceneFuse-3D produces scene assets with clear structure, consistent layout, and fine-grained surface details that closely match the reference top-down image. In contrast, Trellis often generates overly centralized, low-resolution structures and lacks peripheral detail. Hunyuan3D-2 exhibits notable issues with layout distortion and geometry hallucinations despite acceptable textures in isolated parts. TripoSG maintains some compositional structure but frequently introduces repeated objects and ignores the layout evidence within the reference image. LGM can only produce a hollow cube with the scene texture mapped onto its faces. SceneFuse-3D ’s region-wise generation and spatial inpainting pipeline helps it avoid these artifacts while maintaining both global coherence and local fidelity.
These findings confirm that spatial decomposition and prior-guided inpainting are effective principles for lifting single-view image inputs into coherent, high-quality 3D scenes. Additional qualitative comparisons are available in Figure 5 in Appendix A.1.
Table 2: Rendered-view metrics on the reference images. CLIP similarity and FID are computed between each method’s render and the corresponding reference image.
Models CLIP $\uparrow$ FID $\downarrow$
SceneFuse-3D (Ours) 0.8030 258.74
Trellis (Xiang et al., 2024) 0.7760 288.23
Hunyuan3D-2 (Team, 2025) 0.7716 302.26
TripoSG (Li et al., 2025) 0.7203 353.78
LGM (Tang et al., 2024b) 0.7510 353.86
@@ 4.3 Ablation Study
Table 3: Ablation Study Results. Win rates for geometry, layout, and texture show that removing region-based generation or landmark conditioning degrades performance, highlighting the importance of both components in SceneFuse-3D.
@@
SceneFuse-3D (w/o regions) 20.00% 17.00% 13.50% 6.11% 8.48% 6.08%
SceneFuse-3D 80.00% 83.00% 86.50% 92.89% 91.52% 92.92%
SceneFuse-3D (w/o landmarks) 41.50% 46.00% 42.00% 36.90% 44.21% 38.26%
SceneFuse-3D 59.50% 54.00% 58.00% 64.10% 56.79% 61.74%
Figure 4: Qualitative Ablation Results. Left: Reference image. Middle: SceneFuse-3D without landmark conditioning. Right: SceneFuse-3D without region-based generation. Landmark conditioning ensures consistency for foreground objects, especially for objects across regions, while region-based generation preserves overall detail and coherence.
We conduct ablation studies to evaluate the contributions of key components in SceneFuse-3D: the region-based generation strategy and the use of pre-generated landmarks. Both ablation studies still apply the spatial-aware 3D completion. Quantitative results are shown in Table 3, and qualitative comparisons are illustrated in Figure 4.
@@
Removing landmark-aware initialization (depth-only prior) degrades geometry and layout, especially around large foreground structures (e.g., gates/towers). Landmarks act as semantic and geometric anchors across patches: they fix orientation/scale and provide reliable context for masked completion. Without them, region completions drift at boundaries, yielding duplicated façades or misaligned parts across neighboring patches (Fig. 4, middle).
Overall, these ablations show complementary roles: region-wise decomposition keeps generation within the base model’s effective receptive field and preserves local detail, while landmark anchors enforce cross-patch continuity and stabilize global structure—both are necessary for coherent, high-fidelity scenes.
@@ 5 Conclusion
To address the challenge of generating high-quality, coherent 3D scenes from a single image, we proposed SceneFuse-3D, a training-free framework that decomposes scenes into overlapping regions and guides generation with spatial priors. A spatial-aware 3D completion with masked rectified flow preserves local object fidelity while enforcing global coherence. Empirically, SceneFuse-3D outperforms existing methods across geometry, texture, and layout, underscoring the promise of modular, spatially grounded generation for stereotype 3D scene synthesis from minimal input.
@@ Appendix A More Results
A.1 Additional Qualitative Results
Figure 5 presents additional qualitative results comparing SceneFuse-3D with Trellis (Xiang et al., 2024), Hunyuan3D-2 (Team, 2025), and TripoSG (Li et al., 2025) across a diverse set of visual scenes. These examples further demonstrate the robustness and generality of our approach across different architectural styles, spatial layouts, and artistic domains.
SceneFuse-3D consistently produces scene assets that are geometrically detailed, visually coherent, and well-aligned with the input reference images. In contrast, baseline models frequently suffer from artifacts such as repeated structures, layout collapse, or low-resolution textures. These comparisons further highlight the benefits of our region-based generation and spatial-aware inpainting pipeline in generating diverse 3D scenes from a single image without training.
A.2 Additional Baseline
WonderWorld (Yu et al., 2025) enables novel view synthesis of 3D scenes from a single image through text-guided inpainting, which is designed to be effective when a coarse single-view Gaussian representation is initialized and can be refined and enriched across different views. In our task of constructing a consistent 3D scene, we set up the camera trajectory to rotate along the vertical normal to the horizontal plane of the reference images for a comprehensive scene synthesis. However, WonderWorld struggles to generate view-consistent structures and often fails to update the Gaussians effectively due to suboptimal inpainting results based on the existing observations, shown in Fig. 6.
@@ Appendix C Experiment Settings
To evaluate models’ performance under stylistically diverse conditions, we curated a human-verified synthesized image test set with 100 top-down views. Given an example image, we ask ChatGPT-o3 (OpenAI, 2025) to generate image prompts for top-down scene views with these requirements: 1280 × 720 resolution, quasi-orthographic three-quarter (“isometric”) camera, one hero landmark at the image centre, 10–20 surrounding buildings, daylight illumination, and “no far-away object”. Then, we used GPT-4o (Hurst et al., 2024) to generate scene images according to the corresponding prompts.
@@
For GPT-based automatic evaluation, we ask the same questions as the human evaluation and prompt the model to directly return the answer with top-5 token log probability. Then, we extract the token probability of ‘A’ and ‘B’ (0 if not included) as the answer weights. We treat $P(\text{A})$ as a soft vote when our method is option A (and analogously for B). Weighted win rate is computed as $P(\text{win})$ if our model wins, and $1-P(\text{lose})$ otherwise, then averaged across all pairs.
@@ Appendix D Limitation
Although SceneFuse-3D delivers strong scene‐level results, several challenges remain. The pretrained 3D generator we adopt is trained on single‐object imagery; even after region decomposition, the underlying distribution mismatch can lead to patch-level hallucinations—for example, duplicated façades or unrealistic roof shapes. Future work could mitigate this via scene-level fine-tuning or domain adaptation.